# Adversarial Agents: Black-Box Evasion Attacks with Reinforcement Learning

### v1

@@ Abstract
Reinforcement learning (RL) offers powerful techniques for solving complex sequential decision-making tasks from experience. In this paper, we demonstrate how RL can be applied to adversarial machine learning (AML) to develop a new class of attacks that learn to generate adversarial examples: inputs designed to fool machine learning models. Unlike traditional AML methods that craft adversarial examples independently, our RL-based approach retains and exploits past attack experience to improve future attacks. We formulate adversarial example generation as a Markov Decision Process and evaluate RL’s ability to (a) learn effective and efficient attack strategies and (b) compete with state-of-the-art AML. On CIFAR-10, our agent increases the success rate of adversarial examples by 19.4% and decreases the median number of victim model queries per adversarial example by 53.2% from the start to the end of training. In a head-to-head comparison with a state-of-the-art image attack, SquareAttack, our approach enables an adversary to generate adversarial examples with 13.1% more success after 5000 episodes of training. From a security perspective, this work demonstrates a powerful new attack vector that uses RL to attack ML models efficiently and at scale.
@@ 1 Introduction
The advancement of AI has led to an explosion of applications that rely on the decision-making and generation capabilities of ML models. Such models have revolutionized online assistants, commerce, content generation, cybersecurity, entertainment and gaming, to name just a few. At the same time, adversaries wishing to manipulate these applications have developed algorithms that produce inputs that are intended to fool the model, e.g., bypass a content filter. The related field–called adversarial machine learning (AML)–studies algorithms for adversarial examples targeted at victim models (Papernot et al., 2016). Extant AML uses optimization algorithms to minimize the distortion (called perturbations) applied to an input such that the resultant input with distortion is misclassified by the victim ML model (Madry et al., 2019; Carlini & Wagner, 2017). However, the technical community’s understanding of model defenses and adversarial capabilities is at best limited.
Adversarial example generation algorithms rely on gradient-based optimization that operate in isolation from any other information (Madry et al., 2019; Carlini & Wagner, 2017; Andriushchenko et al., 2020). These approaches fail to leverage information gained from previous attacks to improve subsequent attacks on other data. This presents a gap in AML that studies if adversarial examples can be learned, i.e., if the effectiveness and efficiency of the attack improves with experience. If an adversary with restricted access (called black-box access) to a victim model were able to produce adversarial examples at scale (e.g., distributed denial of service attack), this would represent a major new capability for adversaries attacking machine learning.
We posit that modeling an AML adversary as an RL agent will enable attack strategies that become increasingly more efficient and effective over time–which is a capability not currently available in existing work. In this paper, we introduce and evaluate an attack based on RL to generate adversarial examples. Modeled as an RL agent, the adversary learns which perturbations are the most effective at fooling the model. Once a policy is learned, the adversary then uses the policy to generate adversarial examples. Thus, an adversarial agent can attack a model without a costly gradient optimization.
@@
Figure 1: Overview of the reinforcement learning attack with CIFAR-10. The adversary interacts with the victim model by querying images and receiving feedback, iteratively generating adversarial examples while adapting its strategy based on new input data.
@@
We demonstrate that adversarial example generation can be modeled as a Markov Decision Process (MDP). The MDP formulation allows us to readily use RL in such a way that encapsulates the underlying semantics of the attack: input samples and victim model outputs as states, perturbations as actions, and differences in adversarial objectives as rewards. We introduce two RL attacks that reflect the two categories of traditional AML algorithms: RL Max Loss and RL Min Norm. The attacks use an RL algorithm to train the agent, then are used in a policy evaluation setting to craft adversarial examples. The attacks are performed on the CIFAR-10 image classification task against a ResNet50 (He et al., 2015) victim model with the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017). Under this framework, we perform a multi-step evaluation of the agent with respect to (a) learning, (b) tuning and (c) accuracy relative to known approaches.
@@
First, we evaluate if an RL agent can learn adversarial examples. In other words, can RL improve the effectiveness and efficiency of adversarial examples over training? Indeed, both RL Max Loss and RL Min Norm attacks increase average return over training, validating RL’s ability to learn the task. Throughout training, the rate at which adversarial examples are produced increases by 19.4% while the amount of interaction with the victim model decreases by 53.2%, on average. These results empirically show that agents grow stronger over training by producing more adversarial examples efficiently. We show the generation RL-based adversarial examples in Figure 1.
@@
Next, we analyze how the reward and transition hyperparameters in both RL Max Loss and RL Min Norm affect adversarial examples. The $\epsilon$ parameter controls the amount of distortion the RL Max Loss agent is allowed to have on a given input. The $c$ parameter controls the degree to which the RL Min Norm agent gets rewarded for reducing victim model confidence over minimizing distortion. While training the agent increases the performance of both attacks, we show that it is also dependent on the choice of $\epsilon$ and $c$. For our experiments, we choose $\epsilon=0.5$ and $c=1.0$ according to this sensitivity analysis that balances adversarial objectives. An adversary using these attacks must consider them before attacking to meet their desired goal.
@@
Lastly, we evaluate how the trained agent’s ability to craft adversarial examples generalizes to unseen data and how it performs relative to the widely used optimization-based SquareAttack algorithm (Andriushchenko et al., 2020). On an unseen dataset outside of training, the trained agent’s misclassification rate, median queries, and average distortion across adversarial examples fall in distribution of the training adversarial examples. In a black-box comparison with SquareAttack, we compare $5000$ episodes of the RL attacks against state-of-the-art to show that learning adversarial examples enables the RL attacks to generate 13.1% more adversarial examples. These results highlight the efficacy of an adversary learning adversarial examples through RL over existing methods.
2 Background
Evasion attacks in AML aim to craft adversarial examples: inputs that create a disagreement between human and machine, e.g., making a picture that a human would recognize as bus be classified as an ostrich  (Szegedy et al., 2014). In this the adversary attempts to minimize the perturbation applied to an input to reduce the victim model’s confidence on the true label. Crafting adversarial examples balances (a) minimizing the victim model’s confidence on the true label and (b) minimizing the amount of distortion applied. Here, we describe two classes of algorithms that optimize for one while holding the other to a constraint:
[…]
where $\delta$ is the distortion, $x\in\mathbb{R}^{n}$ the input, $y$ the true label, $\epsilon$ the distortion budget, $L$ the loss function between victim model predictions and true label (e.g., cross-entropy loss), and $Z:\mathbb{R}^{n}\to\mathbb{R}^{m}$ the victim model classifier. Max Loss attacks take steps in the direction of maximizing the victim model loss while enforcing budget constraints by projecting onto the $\ell_{p}$-norm budget ball  (Madry et al., 2019; Andriushchenko et al., 2020). Min Norm attacks optimize a soft constraint optimization objective and use a search algorithm to select the multiplier that enforces the misclassification constraint while taking steps that minimize the $\ell_{p}$-norm of distortion  (Carlini & Wagner, 2017).
Referred to in the security domain as a threat model, solving the optimization depends on access to the victim model $Z$. Here, we define our threat model mode with respect to these classes of attacks:
White-Box Attacks This setting is when the adversary is assumed to have full access to the victim model and its parameters. These attacks are stronger because they can readily apply first and second order gradient methods like projected gradient descent (Max Loss) or stochastic gradient descent on a soft constraint optimization (Min Norm) (Madry et al., 2019; Carlini & Wagner, 2017). These attacks help us understand what the worst-case adversary can do to a machine learning model.
Black-Box Attacks This setting assumes only query-based access to the victim model, i.e., only inputs and outputs. Black-box attacks can be further divided into decision-based and score-based: if the output of the model is a classification (i.e., hard-label) or a vector of class probabilities. Despite limited access, these attacks are effective in achieving their goal (Lucas et al., 2023; Sheatsley et al., 2022; Andriushchenko et al., 2020). While minimizing victim model confidence and distortion are the main AML benchmarks, a third benchmark exists in minimizing victim model queries for black-box attacks. SquareAttack is a score-based black-box attack on images that iteratively crafts random square patches and keeps the patch if it increases the victim model loss (Max Loss) (Andriushchenko et al., 2020). In this paper, we consider a score-based black-box threat model where RL can use victim model confidences for a well-shaped reward.
@@ 3 Learning Adversarial Policies
RL enables the adversary to learn from the success and failure of crafting adversarial examples to improve future attacks. Black-box settings involve a feedback loop between the adversary and victim model, allowing us to construct the MDP necessary to use RL. The rest of the section will layout (1) the episodic setting with the MDP, and (2) the attack procedure.
@@ 3.1 MDP Formulation
In the adversary’s MDP, each episode begins with a randomly sampled clean input as part of the initial state. During an episode, the adversary applies perturbations to a select set of features and is rewarded in part by the decrease in victim model confidence on the true label of the selected image. Episodes terminate after the victim model is fooled, meaning that an adversarial example has been crafted (or a threshold of steps is reached, and the process has been deemed to fail). Here, we show an example trajectory with a random agent and randomly sampled clean input from CIFAR-10:
@@
3.1.1 States and Actions
The state space of the attack MDP consists of information available to the adversary. In the black-box setting, this includes the victim model outputs on a given sample $Z(\cdot)$ (i.e., oracle). We define the state representation $s_{t}$ of the attack at a given timestep $t$ as:
@@ 3.1.1 States and Actions
where $x_{t}\in\mathbb{R}^{n}$ is the resulting input after $t$ steps of the environment starting at input sample $x_{0}$, $y$ the ground-truth label of $x_{0}$, and $Z(x_{t})$ the victim model outputs on $x_{t}$. This information allows the agent to learn effective perturbations at different states of the process.
The actions in the environment represent a distortion to be applied to the current input $x_{t}$. Learning fine-grained distortions on every feature of the input becomes difficult on high-dimensional data (e.g., images, text, or network data). We define the action $a_{t}$ at a given timestep $t$ as a vector of $N$ features and associated distortions $\delta$:
$a_{t}=\left((i_{1},\delta_{1}),(i_{2},\delta_{2}),\dots,(i_{N},\delta_{N})\right)$ (3)
where $i_{1},i_{2},...,i_{N}\in\{1,n\}$ are the $N$ features of the input $x_{t}$ selected. Each selected feature is paired with distortions $\delta_{1},\delta_{2},...,\delta_{N}$ with each having magnitude $|\delta_{j}|\leq\theta$ for $j\in\{1,...,N\}$. This allows us to balance the trade-off between many and few distorted features with large or little distortion through $N$ and $\theta$ for effective and efficient policy learning.
3.1.2 Reward and Transition Functions
As highlighted above, we categorize AML objectives into two classes: Max Loss and Min Norm. Our goal is to engineer reward and transition functions so that RL optimizes these objectives. To quantify the victim model’s confidence of the true label, we use:
$f(x,y)=\max\{[Z(x)]_{y}-\max_{i\neq y}[Z(x)]_{i},0\}$ (4)
where $f$ measure the difference between the victim model’s confidence on the true label $y$ and the highest confidence value that is not of the true label $y$. The function is clipped at 00, which means that any iterate $x_{t}$ that is misclassified by the victim model will have a value $f(x,y)=0$. To model one-step differences in victim model confidence and distortion, let us define:
$\Delta_{t+1}f=f(x_{t},y)-f(x_{t+1},y)\quad;\quad\Delta_{t+1}\delta=\|x_{t}-x_{0}\|_{2}-\|x_{t+1}-x_{0}\|_{2}$ (5)
@@ 3.1.2 Reward and Transition Functions
$x_{t+1}=\big{\{}x_{t}^{a_{t}}\textbf{ if }f(x_{t},y)-f(x_{t}^{a_{t}},y)>0\textbf{ otherwise }x_{t}\big{\}}$ (6)
@@
$x_{t+1}=\big{\{}x_{t}^{a_{t}}\textbf{ if }c\cdot\big{[}f(x_{t},y)-f(x_{t}^{a_{t}},y)\big{]}+\big{[}\|x_{t}-x_{0}\|_{2}-\|x_{t}^{a_{t}}-x_{0}\|_{2}\big{]}>0\big{\}}$ (8)
@@
where $c$ weights the importance of reducing the victim model confidence versus minimizing distortion. We experiment with different values of $c$ in the evaluation to determine a value that balances the effectiveness and efficiency of distortions. Like RL Max Loss, a transition involves one query and the distortion proposed by the action is kept if the change in victim model loss and distortion is positive and rewarded accordingly. We denote RL Min Norm transition and reward as $P_{\text{Min Norm}}$ and $R_{\text{Min Norm}}$, respectively.
@@ 3.2 The Reinforcement Learning Attack
We leverage the MDP to launch the RL attack. The attack comes in two forms: (a) attacking while training the agent [^1] Note that in practice, an attacker would train the agent by accessing the model in a subtle way to avoid detection (e.g., by querying the model slowly over time or using known model training data to create a surrogate model to train the agent, (Papernot et al., 2017)). Thereafter, the adversary could use the trained agent to attack the model without restraint. and (b) attacking with a fixed policy. In Figure 2, we detail the attack during the training stage. First, a randomly sampled input and label $(x_{0},y)$ from training dataset $\mathcal{D}$ is used to query the victim model to initialize the start state $s_{0}=(x_{0},y,Z(x_{0}))$. The $(s,a,r,s^{\prime})$ interaction with the victim model proceeds following policy $\pi_{\theta}$ with respect to the attack type RL Max Loss or RL Min Norm. The agent stores interactions and updates the policy according to the specific RL algorithm (on-policy or off-policy). When the episode terminates, the process restarts at start state sampling. After training the agent, the adversary can use the trained adversarial agent to perform the attack in a traditional RL policy evaluation setting. The next section evaluates how well this approach trains adversarial agents that learn better attack strategies.

### v2

@@ Abstract
Attacks on machine learning models have been extensively studied through stateless optimization. In this paper, we demonstrate how a reinforcement learning (RL) agent can learn a new class of attack algorithms that generate adversarial samples. Unlike traditional adversarial machine learning (AML) methods that craft adversarial samples independently, our RL-based approach retains and exploits past attack experience to improve the effectiveness and efficiency of future attacks. We formulate adversarial sample generation as a Markov Decision Process and evaluate RL’s ability to (a) learn effective and efficient attack strategies and (b) compete with state-of-the-art AML. On two image classification benchmarks, our agent increases attack success rate by up to 13.2% and decreases the average number of victim model queries per attack by up to 16.9% from the start to the end of training. In a head-to-head comparison with state-of-the-art image attacks, our approach enables an adversary to generate adversarial samples with 17% more success on unseen inputs post-training. From a security perspective, this work demonstrates a powerful new attack vector that uses RL to train agents that attack ML models efficiently and at scale.
@@ 1 Introduction
The advancement of AI has led to an explosion of applications that rely on the decision-making and generation capabilities of ML models. Such models have revolutionized online assistants, commerce, content generation, cybersecurity, entertainment and gaming, to name just a few. At the same time, adversaries wishing to manipulate these applications have developed algorithms that produce inputs that are intended to fool the model, e.g., bypass a content filter. Adversarial machine learning (AML) studies algorithms for generating adversarial samples targeted at victim models [papernot_limitations_2016, carlini_towards_2017, madry_towards_2017]. Extant AML uses gradient-based optimization to minimize the distortion applied to an input such that the result is misclassified by the victim ML model [chen_zoo_2017, chen_hopskipjumpattack_2020, croce_reliable_2020, vedaldi_square_2020]. However, the technical community’s understanding of model defenses and adversarial capabilities is at best limited.
@@
Figure 1: Overview of the reinforcement learning attack with CIFAR-10. The adversary interacts with the victim model by querying images and receiving feedback, iteratively generating adversarial samples.
Most existing adversarial algorithms are stateless, treating each target input as a one-off optimization problem. This presents a gap in AML, as the adversary is unable to learn and improve its strategy from past experience. For example, consider an advanced persistent threat (APT) agent that uses reinforcement learning (RL) to evade facial recognition systems with restricted query access (i.e., black-box access). This sustained interaction allows the APT agent to learn a generalizable evasion strategy that successfully attacks multiple inputs, extending beyond the limitations of single-sample attacks. Such learned persistence and efficiency would represent a new capability for adversaries attacking machine learning systems.
We posit that modeling the adversary as a reinforcement learning (RL) agent will enable attack strategies that become increasingly more efficient and effective over time, a capability not currently available in existing work. In this paper, we introduce and evaluate an attack based on RL primitives to generate adversarial samples. Modeled as an RL agent, the adversary learns which perturbations are the most effective at fooling the model given an arbitrary input. Once the agent is trained, the adversary then uses the learned policy to generate adversarial samples. Thus, an adversarial agent can attack a model without costly gradient optimization.
@@
We demonstrate that adversarial sample generation can be modeled as a Markov Decision Process (MDP). The MDP formulation allows us to readily use RL in such a way that encapsulates the underlying semantics of the attack: inputs and victim model outputs as states, perturbations as actions, and differences in adversarial objectives as rewards. We introduce two distinct MDPs, RL Max Loss and RL Min Norm, each designed to achieve a different adversarial objective. For both MDPs, we use the Proximal Policy Optimization (PPO) algorithm [schulman_proximal_2017] to train an agent, which is then used in a policy evaluation setting to craft adversarial samples. The attacks are demonstrated on two benchmark image classification tasks against ResNet50 [he_deep_2015], VGG-16 [simonyan_very_2015], and ViT-B/16 [dosovitskiy_image_2021] victim models. Within this framework, we conduct a multi-step evaluation of the agent, assessing its performance across (a) learning capabilities, (b) hyperparameter tuning, and (c) accuracy relative to a standard black-box attack algorithm.
@@
First, we evaluate whether an RL agent can learn strategies that improve black-box evasion attacks. In other words, can RL improve the effectiveness and efficiency of adversarial samples over training? Indeed, both RL Max Loss and RL Min Norm attacks increase discounted return over training, validating RL’s ability to learn the task. Throughout training, the rate at which adversarial samples are produced increases by up to 13.2%, while the amount of interaction with the victim model decreases by up to 16.9%. These results empirically show that agents grow stronger with training by producing more adversarial samples efficiently. We illustrate adversarial agent training with RL on CIFAR-10 samples in Figure 1.
@@
Next, we analyze how MDP hyperparameters in both RL Max Loss and RL Min Norm affect adversarial samples. The $\epsilon$ parameter controls the amount of distortion the RL Max Loss agent is allowed to have on a given input. The $c$ parameter controls the degree to which the RL Min Norm agent gets rewarded for minimizing distortion over reducing victim model confidence. While training the agent increases the performance of both attacks, we show that trained agent performance depends on the choice of $\epsilon$ and $c$. For our experiments, we choose $\epsilon$ and $c$ according to this sensitivity analysis that balances adversarial objectives. An adversary using these attacks must consider them before attacking to meet their desired goal.
@@
Lastly, we evaluate how the trained agent’s ability to craft adversarial samples generalizes to unseen data and how it performs relative to traditional query-based black-box attack algorithms [andriushchenko_square_2020, chen_hopskipjumpattack_2020, ilyas_prior_2019]. On an unseen dataset, the trained agent’s attack success rate, average queries, and average distortion remain consistent with the distribution observed during the training of adversarial samples. In a black-box comparison with the optimization-based baselines, we compare the trained agent’s performance on unseen data to show that using RL to improve black-box evasion attacks enables the adversary to generate up to 17% more adversarial samples with 31% fewer queries. These results highlight the efficacy of an adversary learning from past attack experience through RL over existing methods.
2 Background and Related Work
Evasion attacks in AML aim to craft adversarial samples: inputs, typically with human-imperceptible modifications, that cause a model to make an incorrect prediction [szegedy_intriguing_2014, goodfellow_explaining_2015]. Crafting these samples involves a trade-off between two primary objectives: (a) maximizing the victim model’s prediction error for a given distortion budget (Max Loss), or (b) finding the minimum possible distortion required to cause a misclassification (Min Norm). Formally, given an input x, its true label y, and the victim model classifier Z, these objectives are defined as:
$\text{{Max Loss: }}\arg\max_{\delta}L(Z(x+\delta),y)\text{ s.t. }\|\delta\|_{p}\leq\epsilon$ (1)
$\text{{Min Norm: }}\arg\min_{\delta}\|\delta\|_{p}\text{ s.t. }\arg\max Z(x+\delta)\neq y$
where $\delta$ is the adversarial perturbation, $L$ is a loss function (e.g., cross-entropy), and $\epsilon$ is the maximum distortion budget for a given $\ell_{p}$-norm. The method for solving these optimizations depends on the adversary’s knowledge, or threat model. In a white-box setting, full access to the victim model’s architecture and parameters enables efficient, gradient-based methods like Projected Gradient Descent (PGD) for Max Loss attacks [madry_towards_2017] and the Carlini & Wagner (C&W) attack for Min Norm problems [carlini_towards_2017].
In the more challenging black-box setting, the adversary has only query access to the model’s outputs. These outputs can be hard-label class predictions (decision-based) or class probabilities (score-based). For black-box attacks, minimizing the number of queries becomes a third critical objective alongside maximizing model loss and minimizing distortion. A prominent score-based black-box method used in our evaluation is Square Attack [andriushchenko_square_2020], which solves the max loss problem by iteratively applying random, localized square-shaped updates to the input.
Prior applications of RL to adversarial tasks have focused on the traditional AML setting of optimizing over inputs independently [tsingenopoulos_autoattacker_2019, sarkar_robustness_2023, hore_deep_2025]. We argue that these works understate the true agentic capability of RL in this domain: learning attack strategies that improve with experience and generalize to new data. By drawing a parallel between the objective of AML optimizations and the RL process, our work introduces a generalizable framework built on this paradigm of learning from past attacks. To demonstrate the practicality of this approach, we conduct a utility analysis on data outside the RL agent’s training environment and present a direct comparison of attack success and query-efficiency against established black-box attack methods.
@@ 3 Learning Adversarial Policies
RL enables the adversary to learn from the success and failure of crafting adversarial samples. Black-box settings involve a feedback loop between the adversary and victim model, allowing us to construct the MDP necessary to use RL. The rest of the section will lay out (1) the episodic setting with the MDP, and (2) the attack procedure.
@@ 3.1 MDP Formulation
In the adversary’s MDP, each episode begins with a randomly sampled clean input as part of the initial state. During an episode, the adversary applies perturbations to a select set of features and is rewarded in part by the decrease in victim model confidence on the true label of the selected image. Episodes terminate after the victim model is fooled, meaning that an adversarial sample has been crafted (or a threshold of steps is reached, and the process has been deemed to fail). Here, we show an episode with a random agent and randomly sampled input from CIFAR-10 [krizhevsky_learning_nodate]:
@@
States and Actions. The state space of the attack MDP consists of information available to the adversary. In the black-box setting, this includes the victim model outputs on a given sample $Z(\cdot)$ (i.e., oracle). We define the state representation $s_{t}$ of the attack at a given timestep $t$ as:
@@
where $x_{t}\in\mathbb{R}^{n}$ is the resulting input after $t$ steps of the environment starting at input sample $x_{0}$, $y$ the ground-truth label of $x_{0}$, and $Z(x_{t})$ the victim model outputs on $x_{t}$. This information allows the agent to learn effective perturbations at different states of the process. The actions in the environment represent a distortion to be applied to the current input $x_{t}$. Learning fine-grained distortions on every feature of the input becomes difficult on high-dimensional data (e.g., images, text, or network data). To create a more tractable learning problem, we simplify the action space by having the agent select a small subset of $N$ features to perturb. This dimensionality reduction is critical for query efficiency in black-box settings, as searching the full high-dimensional input space is often infeasible [ilyas_prior_2019]. We define the action $a_{t}$ at a given timestep $t$ as a set of $N$ feature-perturbation pairs:
$a_{t}=\left\{(i_{1},\delta_{1}),(i_{2},\delta_{2}),\dots,(i_{N},\delta_{N})\right\}$ (3)
where $i_{1},...,i_{N}\in\{1,n\}$ are the indices of the $N$ input features selected for modification. Each selected feature $i_{j}$ paired with a distortion $\delta_{j}$, with each having magnitude $|\delta_{j}|\leq\theta$ for $j\in\{1,...,N\}$. This allows us to balance the trade-off between many and few distorted features with large or little distortion through $N$ and $\theta$ for effective and efficient policy learning.
Reward and Transition Functions. As highlighted above, we categorize AML objectives into two classes: Max Loss and Min Norm. Our goal is to shape reward and transition functions so that RL optimizes these objectives. To quantify the victim model’s confidence of the true label, we use:
$f(x,y)=\log([Z(x)]_{y})$ (4)
where $f$ measures the log probability of the victim model’s confidence of $x$ belonging to class $y$ (i.e., negative cross-entropy loss). To model one-step differences in victim model confidence and distortion, let us define:
$\Delta_{t+1}f=f(x_{t},y)-f(x_{t+1},y)\quad$ (5)
$\Delta_{t+1}\delta=\|x_{t}-x_{0}\|_{2}-\|x_{t+1}-x_{0}\|_{2}$
@@
$x_{t+1}=\begin{cases}x_{t}^{a_{t}}&\textbf{if }f(x_{t},y)-f(x_{t}^{a_{t}},y)>0\\ x_{t}&\textbf{otherwise}\end{cases}$ (6)
@@
$x_{t+1}=\begin{cases}x_{t}^{a_{t}}&\textbf{if }\begin{aligned} &c\cdot\big[\|x_{t}-x_{0}\|_{2}-\|x_{t}^{a_{t}}-x_{0}\|_{2}\big]\\ &+\big[f(x_{t},y)-f(x_{t}^{a_{t}},y)\big]>0\end{aligned}\\ \\ x_{t}&\textbf{otherwise}\end{cases}$ (8)
@@
where $c$ weights the importance of minimizing distortion over reducing the victim model confidence. We experiment with different values of $c$ in the evaluation to determine a value that balances the effectiveness and efficiency of distortions. Like RL Max Loss, a transition involves one query and the distortion proposed by the action is kept if the change in victim model loss and distortion is positive and rewarded accordingly. We denote RL Min Norm transition and reward as $P_{\text{Min Norm}}$ and $R_{\text{Min Norm}}$, respectively.
@@ 3.2 The Reinforcement Learning Attack
We leverage these MDPs to launch the RL attack. The attack comes in two forms: (a) attacking while training the agent [^1] Note that in practice, an attacker would train the agent by accessing the model in a subtle way to avoid detection (e.g., by querying the model slowly over time or using known model training data to create a surrogate model to train the agent, [papernot_practical_2017]). Thereafter, the adversary could use the trained agent to attack the model without restraint. and (b) attacking with a fixed policy. In Figure 2, we detail the attack during the training stage. First, a randomly sampled input and label $(x_{0},y)$ from training dataset $\mathcal{D}$ is used to query the victim model to initialize the start state $s_{0}=(x_{0},y,Z(x_{0}))$. The $(s,a,r,s^{\prime})$ interaction with the victim model proceeds following policy $\pi_{\theta}$ with respect to the attack type RL Max Loss or RL Min Norm. The agent stores interactions and updates the policy according to the specific RL algorithm (on-policy or off-policy). When the episode terminates, the process restarts at start state sampling. After training the agent, the adversary can use the trained adversarial agent to perform the attack in a traditional RL policy evaluation setting. The next section evaluates how well this approach trains adversarial agents that learn better attack strategies.